State 6. The GSS 2014 he GSS 2014 asked respondents to identify what was important for truly being merican. When asked How
Answer #1

ANSWER:

Ert = Z1 = IH : H H, : Atleast two 4; differ х X2 X3 9. 16 4 4 4 4 4 4 4 4 3 3 4 16 4 4 16 4 ΣΧ ΣΧ Σ ΣΧ3 Σ ΣΧ) 20 62 11 19 15k = 3 (Number of groups or treatment Conditions) n = 20 (Number of data points.) %3D (ΣΧ+. +Σ x)2 Ση SSTotal = (Exỉ +...+Ex3)Significance level = 0.01 MSAmong MSwithin %3D 6.128571/2 10.071429/17 SSAmong /k – 1 SSwithin/n - 3 Test statistic, F 3.0642ANOVA Source df MS 6.128571 3.064286 5.172 Among 2 Within 10.071429 0.592437 17 Total 19 16.2Degrees of freedom for Numerator, d fN = k – 1 = 2 Degrees of freedom for Denominator, dfp = n- k = 17 To get F-critical-valu